OpenStudy (anonymous):

the distribution center receives shipments of a product from three different factories in the quantities of 50, 40, and 25 three times a product is selected at random, each time without replacement find the probability of (A) all three products came from the third factory (B) none of the three products came from the third factory

OpenStudy (anonymous):

You have the chance of 1/3 that the first will be picked from the third. 1/3 second 1/3 for third. 1/3 * 1/3 * 1/3 = 1/27

OpenStudy (anonymous):

now it cant come from the third so 2/3 * 2/3 * 2/3 = 8/27

OpenStudy (anonymous):

If you have trouble seeing how it can work, you can create a probability tree. Works great.

OpenStudy (anonymous):



OpenStudy (anonymous):

It's incomplete but you should get the idea

OpenStudy (anonymous):

what is the probability that all three products came from the same factory? round to the nearest 1000

OpenStudy (anonymous):

1/27 = 0.03703703703

OpenStudy (anonymous):

Remember... when you're rounding in decimals, you start by tenth

OpenStudy (anonymous):

So 0.037

OpenStudy (anonymous):

where are you getting the 27? its 25 and it does not replace it

OpenStudy (anonymous):

oh. i misread your entire question.

OpenStudy (anonymous):

misread my question - thanks clown

OpenStudy (anonymous):

I did try, I don't do mistakes on purpose. Perhaps a "Nice try" would be something more convenient.

OpenStudy (mathmate):

have a feeling the question has a line (or part of a line) skipped. Could you reread the question for typo? The question sounds misleading, and may actually mean that all products are then mixed, and out of the whole batch, 3 items are picked. Kindly check for missing information or typo.

OpenStudy (goformit100):

Hello, and A Warm Welcome to Open Study!